OpenStudy (anonymous):

Decompose -2x-23/2x^2-9x-5 into partial fractions

OpenStudy (anonymous):

\[\frac{ -2x-23 }{ 2x ^{2}-9x-5 }\]

OpenStudy (anonymous):

i got \[\frac{ -3 }{ 2x+1 }+\frac{ 4 }{ x-5 }\]